OpenStudy (anonymous):

Solve for x. Round to the nearest ten-thousandth. 0.12x = 10-3. please just give me first step to start

OpenStudy (anonymous):

it 0.1^2x=10^-3

OpenStudy (lilg132):

0.12x = 7 To get x on its own you divide it by the number next to it and you also divide the other side by 0.12 0.12x / 0.12 = 7 / 0.12 x = 58.33

OpenStudy (anonymous):

where are you getting the 7

OpenStudy (anonymous):